# M3-VRD: Multimodal Multi-task Multi-teacher Visually-Rich Form Document Understanding

### v2

@@ Abstract
This paper presents a groundbreaking multimodal, multi-task, multi-teacher joint-grained knowledge distillation model for visually-rich form document understanding. The model is designed to leverage insights from both fine-grained and coarse-grained levels by facilitating a nuanced correlation between token and entity representations, addressing the complexities inherent in form documents. Additionally, we introduce new inter-grained and cross-grained loss functions to further refine diverse multi-teacher knowledge distillation transfer process, presenting distribution gaps and a harmonised understanding of form documents. Through a comprehensive evaluation across publicly available form document understanding datasets, our proposed model consistently outperforms existing baselines, showcasing its efficacy in handling the intricate structures and content of visually complex form documents.
@@ 1 Introduction
Understanding and extracting structural information from Visually-Rich Documents (VRDs), such as academic papers Zhong et al. (2019); Ding et al. (2023b), receipts Park et al. (2019), and forms Jaume et al. (2019); Ding et al. (2023a), holds immense value for Natural Language Processing (NLP) tasks, particularly in information extraction and retrieval. While significant progress has been made in solving various VRD benchmark challenges, including layout analysis and table structure recognition, the task of form document understanding remains notably challenging. This complexity of the form document understanding arises from two main factors: 1) the involvement of two distinct authors in a form and 2) the integration of diverse visual cues. Firstly, forms mainly involve two primary authors: form designers and users. Form designers create a structured form to collect necessary information as a user interface. Unfortunately, the form layouts, designed to collect varied information, often lead to complex logical relationships, causing confusion for form users and heightening the challenges in form document understanding. Secondly, diverse authors in forms may encounter a combination of different document natures, such as digital, printed, or handwritten forms. Users may submit forms in various formats, introducing noise such as low resolution, uneven scanning, and unclear handwriting. Traditional document understanding models do not account for the diverse carriers of document versions and their associated noises, exacerbating challenges in understanding form structures and their components. Most VRD understanding models inherently hold implicit multimodal document structure analysis (Vision and Text understanding) knowledge either at fine-grained Huang et al. (2022); Wang et al. (2022) or coarse-grained Tan and Bansal (2019); Li et al. (2019) levels. The fine-grained only models mainly focus on learning detailed logical layout arrangement, which cannot handle complex relationships of multimodal components, while the coarse-grained models tend to omit significant words or phrases. Hence, we introduce a novel joint-grained document understanding approach with multimodal multi-teacher knowledge distillation. This method leverages knowledge from various task-based teachers (pre-trained models) throughout the training process, intending to create more inclusive and representative multi- and joint-grained document representations.
Our contributions can be summarised as follows: 1) We present a groundbreaking multimodal, multi-task, multi-teacher joint-grained knowledge distillation model designed explicitly to understand visually-rich form documents. 2) The proposed model outperforms across publicly available form document datasets. 3) To the best of our knowledge, this research marks the first in adopting multi-task knowledge distillation, focusing on incorporating multimodal form document components.
@@ 2 Related Works
Multimodal approaches: They leverage both the recognised text and the spatial relations (denoted by T and S) between document components (see, for instance, LiLT Wang et al. (2022) and BROS Hong et al. (2022)). The main goal is to complement raw content understanding with layout information. Expanding upon this multimodal framework, models such as LayoutLMv3 Huang et al. (2022), DocFormerv2 Appalaraju et al. (2023), Fast-StrucText Zhai et al. (2023), and, FormNetV2 Lee et al. (2023a) integrate the visual modality with text and layout information. These approaches are capable of capturing nuances in the document content hidden in prior works. To leverage multimodal relations, these models are typically pre-trained in a multi-task fashion, exploiting a curated set of token- or word-based pre-training tasks, such as masking or alignment.
Our approach aligns with the multimodal model paradigm, distinguishing itself by eschewing generic pre-training tasks reliant on masking, alignment, or NTP. Instead, it leverages the direct extraction of knowledge from multiple teachers, each trained on downstream datasets, encompassing both entity and token levels of analysis with the proposed inter-grained and cross-grained losses. This enriches the depth of understanding in visual documents, capturing intricate relationships and semantic structures beyond individual tokens.
Figure 1: Multimodal Multi-task Multi-Teacher Joint-Grained Document Understanding Framework. Each section is aligned with the specific colours, Green: Section 3.2.1, Blue: Section 3.2.2, Orange: Section 3.3
@@ 3.3 Multimodal Multi-task Multi-Teacher Knowledge Distillation
This section introduces the multi-loss strategy to enhance inter-grained and cross-grained knowledge exchange, ensuring a more nuanced and effective integration of insights from fine-grained and coarse-grained representations. The accompanying multi-loss ablation study (Section 5.3) aims to optimise the synergies between multiple teacher models, thereby contributing to a more robust and comprehensive joint-grained learning process.
@@ 3.3.1 Task-oriented Cross Entropy Loss
3.3.2 Inter-Grained Loss Functions
@@
The introduction of these inter-grained loss functions, including the similarity loss and the distilling loss, contributes to mitigating distribution gaps and fostering a synchronised understanding of the form across various levels of granularity.
@@ 4 Evaluation Setup
FUNSD Jaume et al. (2019) comprises 199 noisy scanned documents from various domains including marketing, advertising, and science reports, related to US tobacco firms. It is split into train and test sets (149/50 documents), and each document is presented in either printed or handwritten format with low resolutions. Our evaluation focuses on the semantic-entity labeling task that identifies four predefined labels (i.e., question, answer, header, and other) based on input text content.
@@ 5.1 Overall Performance
Table 2: Overall Performance of Model and Configurations on FormNLU printed P and handwritten H. The full form of acronyms can be found in Section 5.1. The best is in bold. The second best is underlined.
@@
To thoroughly distil joint-grained knowledge from multiple teachers, we introduced multiple loss functions encompassing Multiple auxiliary tasks. These functions capture teacher knowledge from inter-grained and cross-grained perspectives, generating representative token embeddings. Typically, using either inter-grained or coarse-grained loss individually leads to better performance than the best baselines across various test sets. Inter-grained Similarity (Sim) and Distilling (Distil) loss consistently achieve higher F1 scores in nearly all test sets. Moreover, cross-grained Triplet and alignment (Align) losses outperform the best baseline on the FormNLU (P) or (H). This highlights the effectiveness of the proposed multi-task learning approach in enhancing token representations by integrating knowledge from joint-grained multi-teachers. Inter-grained loss functions exhibit higher robustness on both datasets, whereas cross-grained loss functions only perform well on FormNLU. This difference may stem from the FUNSD being sourced from multiple origins, whereas FormNLU is a single-source dataset. Coarse-grained loss functions may excel on single-source documents by capturing more prevalent knowledge but might introduce noise when applied to multiple sources. Also, the model demonstrates its most competitive performance by integrating all proposed loss functions (+Sim+Distil+Triplet+Align). This highlights how the proposed inter-grained and cross-grained loss functions enhance multi-teacher knowledge distillation in form understanding tasks [^7] More loss combination analysis is in Section 5.3.
@@ 5.3 Effect of Loss Functions
To comprehensively investigate the impact of different loss functions and their combinations, we present the performance of various combinations in Table 4. While employing inter-grained loss individually often proves more effective than using cross-grained loss alone, combining the two losses can enhance knowledge distillation from joint-grained multi-teachers. For instance, concurrently employing distilling(Distil) and Triplet loss improved accuracy from 97.72% to 98.35%. Notably, stacking all proposed loss functions resulted in the best or second-best performance across all test sets, showcasing their effectiveness in distilling knowledge from multi-teacher to student models for generating more representative representations. Even though cross-grained Triplet and Alignment losses were ineffective individually, when combined with inter-grained loss, they significantly improved knowledge distillation effectiveness.
@@ Appendix B Breakdown Result Analysis
As shown in Table 7, for the FUNSD dataset, we could find all Joint-Grained(JG-) frameworks can have a delicate performance on recognising Question and Answer, but decreased in Header classification. This might result from the limited number of Headers in the FUNSD, leading to inadequate learning of the fine-grained and coarse-grained Header information. Multi-task-oriented inter-grained and coarse-grained functions can increase the performance of Question recognition by boosting the knowledge distilling from joint-grained multi-teachers. Especially, inter-grained knowledge distillation methods can achieve around 1% higher than LayoutLMv3. The FUNSD dataset cannot illustrate the benefits of cross-grained loss functions well.
@@
For FormNLU printed and handwritten sets, the joint-grained framework and proposed loss functions can effectively improve Section (Sec) and Title recognition. As the Title, Section and Form_key (F_K) are normally located at similar positions for single-source forms, this may demonstrate both joint-grained framework and multi-task loss function could distil knowledge. Additionally, baseline models are not good at recognising table keys and values, especially handwritten sets. As we use the layoutLMv3 in the joint-grained framework, the performance of recognising table-related tokens is not good for the joint-learning framework. After integrating multiple teachers, the performance has increased from 91.97% to 97.35% on the printed set. The proposed multi-task loss functions may achieve a higher performance of 97.96%. Significant improvements can also be observed across two test sets across all table-related targets. This illustrates that the joint-grained multi-teacher framework can effectively tackle the limitation of one teacher to generate more comprehensive token representations, and the inter-grained and cross-grained loss could boost the effective knowledge exchange to make the generalisation and robustness of the entire framework.

### v3

# 3MVRD: Multimodal Multi-task Multi-teacher Visually-Rich Form Document Understanding

@@ Abstract
This paper presents a groundbreaking multimodal, multi-task, multi-teacher joint-grained knowledge distillation model for visually-rich form document understanding. The model is designed to leverage insights from both fine-grained and coarse-grained levels by facilitating a nuanced correlation between token and entity representations, addressing the complexities inherent in form documents. Additionally, we introduce new intra-grained and cross-grained loss functions to further refine diverse multi-teacher knowledge distillation transfer process, presenting distribution gaps and a harmonised understanding of form documents. Through a comprehensive evaluation across publicly available form document understanding datasets, our proposed model consistently outperforms existing baselines, showcasing its efficacy in handling the intricate structures and content of visually complex form documents [^1] Code: https://github.com/adlnlp/3mvrd.
@@ 1 Introduction
Understanding and extracting structural information from Visually-Rich Documents (VRDs), such as academic papers Zhong et al. (2019); Ding et al. (2023b), receipts Park et al. (2019), and forms Jaume et al. (2019); Ding et al. (2023a), holds immense value for Natural Language Processing (NLP) tasks, particularly in information extraction and retrieval. While significant progress has been made in solving various VRD benchmark challenges, including layout analysis and table structure recognition, the task of form document understanding remains notably challenging. This complexity of the form document understanding arises from two main factors: 1) the involvement of two distinct authors in a form and 2) the integration of diverse visual cues. Firstly, forms mainly involve two primary authors: form designers and users. Form designers create a structured form to collect necessary information as a user interface. Unfortunately, the form layouts, designed to collect varied information, often lead to complex logical relationships, causing confusion for form users and heightening the challenges in form document understanding. Secondly, diverse authors in forms may encounter a combination of different document natures, such as digital, printed, or handwritten forms. Users may submit forms in various formats, introducing noise such as low resolution, uneven scanning, and unclear handwriting. Traditional document understanding models do not account for the diverse carriers of document versions and their associated noises, exacerbating challenges in understanding form structures and their components. Most VRD understanding models inherently hold implicit multimodal document structure analysis (Vision and Text understanding) knowledge either at fine-grained Huang et al. (2022); Wang et al. (2022) or coarse-grained Tan and Bansal (2019); Li et al. (2019) levels. The fine-grained only models mainly focus on learning detailed logical layout arrangement, which cannot handle complex relationships of multimodal components, while the coarse-grained models tend to omit significant words or phrases. Hence, we introduce a novel joint-grained document understanding approach with multimodal multi-teacher knowledge distillation. It leverages knowledge from various task-based teachers throughout the training process, intending to create more inclusive and representative multi- and joint-grained document representations.
Our contributions are summarised as follows: 1) We present a groundbreaking multimodal, multi-task, multi-teacher joint-grained knowledge distillation model designed explicitly to understand visually-rich form documents. 2) Our model outperforms publicly available form document datasets. 3) This research marks the first in adopting multi-task knowledge distillation, focusing on incorporating multimodal form document components.
@@ 2 Related Works
Multimodal approaches: They leverage both the recognised text and the spatial relations (denoted by T and S) between document components (e.g., LiLT Wang et al. (2022) and BROS Hong et al. (2022)). The main goal is to complement raw content understanding with layout information. Expanding upon this multimodal framework, models such as LayoutLMv3 Huang et al. (2022), DocFormerv2 Appalaraju et al. (2023), Fast-StrucText Zhai et al. (2023), and, FormNetV2 Lee et al. (2023a) integrate the visual modality with text and layout information. These approaches are capable of capturing nuances in the document content hidden in prior works. To leverage multimodal relations, these models are typically pre-trained in a multi-task fashion, exploiting a curated set of token- or word-based pre-training tasks, such as masking or alignment.
Our approach aligns with the multimodal model paradigm, distinguishing itself by eschewing generic pre-training tasks reliant on masking, alignment, or NTP. Instead, it leverages the direct extraction of knowledge from multiple teachers, each trained on downstream datasets, encompassing both entity and token levels of analysis with the proposed intra-grained and cross-grained losses. This enriches the depth of understanding in visual documents, capturing intricate relationships and semantic structures beyond individual tokens.
Figure 1: Multimodal Multi-task Multi-teacher Visually-Rich Form Document Understanding (3MVRD). Each section is aligned with the specific colours, Green: Section 3.2.1, Blue: Section 3.2.2, Orange: Section 3.3
@@ 3.3 Multimodal Multi-task Multi-Teacher Knowledge Distillation
This section introduces the multi-loss strategy to enhance intra-grained and cross-grained knowledge exchange, ensuring a more nuanced and effective integration of insights from fine-grained and coarse-grained representations. The accompanying multi-loss ablation study (Section 5.3) aims to optimise the synergies between multiple teacher models, thereby contributing to a more robust and comprehensive joint-grained learning process.
@@ 3.3.1 Task-oriented Cross Entropy Loss
3.3.2 Intra-Grained Loss Functions
@@
The introduction of these intra-grained loss functions, including the similarity loss and the distilling loss, contributes to mitigating distribution gaps and fostering a synchronised understanding of the form across various levels of granularity.
@@ 4 Evaluation Setup
FUNSD Jaume et al. (2019) comprises 199 noisy scanned documents from various domains, including marketing, advertising, and science reports related to US tobacco firms. It is split into train and test sets (149/50 documents), and each document is presented in either printed or handwritten format with low resolutions. Our evaluation focuses on the semantic-entity labeling task that identifies four predefined labels (i.e., question, answer, header, and other) based on input text content.
@@ 5.1 Overall Performance
Table 2: Overall performance with configurations on FormNLU printed P and handwritten H. The full form of acronyms can be found in Section 5.1. The best is in bold. The best teacher model (baseline) is underlined.
@@
To thoroughly distil joint-grained knowledge from multiple teachers, we introduced multiple loss functions encompassing Multiple auxiliary tasks. These functions capture teacher knowledge from intra-grained and cross-grained perspectives, generating representative token embeddings. Typically, using either intra-grained or coarse-grained loss individually leads to better performance than the best baselines across various test sets. Intra-grained Similarity (Sim) and Distilling (Distil) loss consistently achieve higher F1 scores in nearly all test sets. Moreover, cross-grained Triplet and alignment (Align) losses outperform the best baseline on the FormNLU (P) or (H). This highlights the effectiveness of the proposed multi-task learning approach in enhancing token representations by integrating knowledge from joint-grained multi-teachers. Intra-grained loss functions exhibit higher robustness on both datasets, whereas cross-grained loss functions only perform well on FormNLU. This difference may stem from the FUNSD being sourced from multiple origins, whereas FormNLU is a single-source dataset. Coarse-grained loss functions may excel on single-source documents by capturing more prevalent knowledge but might introduce noise when applied to multiple sources. Also, the model demonstrates its most competitive performance by integrating all proposed loss functions (+Sim+Distil+Triplet+Align). This highlights how the proposed intra-grained and cross-grained loss functions enhance multi-teacher knowledge distillation in form understanding tasks [^8] More loss combination analysis is in Section 5.3.
@@ 5.3 Effect of Loss Functions
To comprehensively investigate the impact of different loss functions and their combinations, we present the performance of various combinations in Table 4. While employing intra-grained loss individually often proves more effective than using cross-grained loss alone, combining the two losses can enhance knowledge distillation from joint-grained multi-teachers. For instance, concurrently employing distilling(Distil) and Triplet loss improved accuracy from 97.72% to 98.35%. Notably, stacking all proposed loss functions resulted in the best or second-best performance across all test sets, showcasing their effectiveness in distilling knowledge from multi-teacher to student models for generating more representative representations. Even though cross-grained Triplet and Alignment losses were ineffective individually, when combined with intra-grained loss, they significantly improved knowledge distillation effectiveness.
@@ Appendix B Breakdown Result Analysis
As shown in Table 7, for the FUNSD dataset, we could find all Joint-Grained(JG-) frameworks can have a delicate performance on recognising Question and Answer, but decreased in Header classification. This might result from the limited number of Headers in the FUNSD, leading to inadequate learning of the fine-grained and coarse-grained Header information. Multi-task-oriented intra-grained and coarse-grained functions can increase the performance of Question recognition by boosting the knowledge distilling from joint-grained multi-teachers. Especially, intra-grained knowledge distillation methods can achieve around 1% higher than LayoutLMv3. The FUNSD dataset cannot illustrate the benefits of cross-grained loss functions well.
@@
For FormNLU printed and handwritten sets, the joint-grained framework and proposed loss functions can effectively improve Section (Sec) and Title recognition. As the Title, Section and Form_key (F_K) are normally located at similar positions for single-source forms, this may demonstrate both joint-grained framework and multi-task loss function could distil knowledge. Additionally, baseline models are not good at recognising table keys and values, especially handwritten sets. As we use the layoutLMv3 in the joint-grained framework, the performance of recognising table-related tokens is not good for the joint-learning framework. After integrating multiple teachers, the performance has increased from 91.97% to 97.35% on the printed set. The proposed multi-task loss functions may achieve a higher performance of 97.96%. Significant improvements can also be observed across two test sets across all table-related targets. This illustrates that the joint-grained multi-teacher framework can effectively tackle the limitation of one teacher to generate more comprehensive token representations, and the intra-grained and cross-grained loss could boost the effective knowledge exchange to make the generalisation and robustness of the entire framework.